wo people are pulling a table (of mass 3 kg) on a hard surface. One person is pulling with
horizontal force of 10 N to the left, the other with a horizontal force of 20 N at -30 ° to the
horizontal. What is the net force on the table?
1
Expert's answer
2020-11-05T10:38:57-0500

Let "F_1=10" N, "F_2=20" N, "\alpha =-30°."

So the net force is "F=F_1+F_2cos\alpha=27.3" N.